OpenStudy (anonymous):

If y varies inversely as x, and y=2/5 when x=2, find y when x=1

Nnesha (nnesha):

equation of inverse variation is \[\huge\rm y = \frac{ k }{ x }\] replace x and y by their value solve for k then plug in k value to find y

OpenStudy (anonymous):

So... K would be .80 ... right?

Nnesha (nnesha):

k =80 hmm how did you get it?

OpenStudy (anonymous):

Well, I converted 2/5 to a decimal, and then multiplied it by 2, which is what x was.

Nnesha (nnesha):

do u mean 0.8 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

It is 4/5.

Nnesha (nnesha):

alright now y = k/x replace k by .8 and x =1 find y

Nnesha (nnesha):

yes that's right

OpenStudy (anonymous):

Do you mind helping me with one more? @Nnesha

Nnesha (nnesha):

i'll try :-)